The force F = i + 3j + 2k acts at the point (1, 1, 1). (a) Find the torque of the force about the point (2, −1, 5). Careful! The vector r goes from (2, −1, 5) to (1, 1, 1)


1
Expert's answer
2022-05-30T15:17:33-0400
"\vec r=(1-2)i+(1-(-1))j+(1-5)k\\
=-i+2j-4k"

The torque

"\vec \tau=\vec r\times \vec F=\begin{vmatrix}
 i & j & k \\
 x & y & z\\
F_x & F_y & F_z
\end{vmatrix}"

"=\begin{vmatrix}
 i & j & k \\
 -1 & 2 & -4\\
1 & 3 & 2
\end{vmatrix}=16i-2j-5k"